OpenStudy (anonymous):

Simplify the complex fraction: [x+(x/2)] over (divided by) [x-(x^2/2)]. I really need help... I'll buy you a pepsi if you get a correct answer for me :)

OpenStudy (mertsj):

\[\frac{x+\frac{x}{2}}{x-\frac{x^2}{2}}\]

OpenStudy (mertsj):

Multiply each term in the numerator and denominator by 2

OpenStudy (mertsj):

\[\frac{2x+x}{2x-x^2}=\frac{3x}{2x-x^2}=\frac{3x}{x(2-x)}=\frac{3}{2-x}\]